# PIP-Net: Pedestrian Intention Prediction in the Wild

### v2

@@ Abstract
Accurate pedestrian intention prediction (PIP) by Autonomous Vehicles (AVs) is one of the current research challenges in this field. In this article, we introduce PIP-Net, a novel framework designed to predict pedestrian crossing intentions by AVs in real-world urban scenarios. We offer two variants of PIP-Net designed for different camera mounts and setups. Leveraging both kinematic data and spatial features from the driving scene, the proposed model employs a recurrent and temporal attention-based solution, outperforming state-of-the-art performance. To enhance the visual representation of road users and their proximity to the ego vehicle, we introduce a categorical depth feature map, combined with a local motion flow feature, providing rich insights into the scene dynamics. Additionally, we explore the impact of expanding the camera’s field of view, from one to three cameras surrounding the ego vehicle, leading to enhancement in the model’s contextual perception. Depending on the traffic scenario and road environment, the model excels in predicting pedestrian crossing intentions up to 4 seconds in advance which is a breakthrough in current research studies in pedestrian intention prediction. Finally, for the first time, we present the Urban-PIP dataset, a customised pedestrian intention prediction dataset, with multi-camera annotations in real-world automated driving scenarios.
@@
Autonomous vehicles, pedestrian crossing behaviour, pedestrian intention prediction, computer vision, deep neural networks.
@@ I Introduction
Pedestrians are the most vulnerable road users and face a high risk of fatal accidents [1]. Ensuring pedestrian safety in automated driving, particularly in mixed AV-pedestrian traffic scenarios, heavily relies on the AV’s capability in “pedestrian intention prediction (PIP)” [2]. A PIP system determines if a pedestrian is likely to cross the road shortly (within the next few seconds). This study aims to investigate the critical visual clues that pedestrians exhibit when they intend to cross the road, and then provide a model which predicts crossing behaviour a few seconds in advance.
Anticipating pedestrian crossing behaviour is a difficult task due to various environmental factors that affect human intention [3, 4]. Even in the simulated scenarios in which the majority of parameters are under control, crossing prediction is a challenging endeavour [5]. Factors like interactions with other pedestrians, traffic signs, road congestion, and vehicle speed can influence pedestrians’ tendency to cross the road in front of AVs [6].
Computer vision plays a crucial role in enabling AVs to perceive their surrounding environment by analysing the visual data captured via multiple sensors, such as cameras, LiDAR, Radar, etc. Learning-based models, in particular deep neural networks (DNNs), have shown remarkable success in various computer vision tasks, including scene understanding, semantic segmentation [7], road users classification, localisation [8], and motion prediction [9]. Figure 1 illustrates some of the perceivable factors such as depth, pedestrian pose, and surrounding objects that an autonomous vehicle should consider to interpret the scene and estimate the pedestrians’ intention. DNNs are particularly effective at learning complex patterns and features from visual data, making them a natural fit for tasks that involve analysing images or videos to comprehend pedestrian behaviour [10, 11, 12]. They also offer significant capabilities in multi-modal integration by providing a neural-based mechanism to process and fuse all the perceived information from diverse sensors. This integration may enhance the overall understanding of the environment and help to make more accurate and safer decisions [13].
@@
Several datasets, such as JAAD [14], PIE [15], and STIP [16], use onboard camera recordings and their data are publicly released for the study of pedestrians’ behaviour before and during road crossing. However, most of the current research works suffer from supplying a multi-camera in-cabin setup to leverage the benefits of sensor fusion and multi-modal perception. In addition to the above-mentioned datasets, some baseline approaches [17] have been also established for analysing the visual cues and signals that pedestrians emit through their body language and positioning. The approaches highlight the benefits of combining these features with contextual information. Contextual information may include factors such as the road’s location, the time of day, weather conditions, the presence of traffic signals or crosswalks, the type of road (urban, suburban, rural), and the position and behaviour of other vehicles near the scene [18]. To the best of our knowledge, no extensive research has been conducted to understand and interpret such contextual details and their effects on pedestrian’s decision-making.
In this study, we propose a customised DNN-based framework, called “PIP-Net” that takes various features of pedestrians, the environment, and the ego-vehicle state into account, to learn the context of a crossing scenario and consequently predict the intention of pedestrians in real-world AV urban driving scenarios. The main contributions of this research are highlighted as follows:
@@
  • Finally, we evaluate the effectiveness of the proposed model on the widely utilised PIE dataset and the introduced Urban-PIP dataset, outperforming the state-of-the-art (SOTA) for predicting pedestrian actions [11].
@@ II Related Works
Recently, pedestrian crossing intention prediction research has surged and gained significant attention within the autonomous driving research community [20, 10, 11, 12]. Most current methods mainly address the problem by taking two aspects into account: Discovering influential factors and features for interpreting road users’ interactions [5, 6, 4], and designing the analytic model to predict the pedestrians’ crossing intention [18, 14, 15, 16, 17]. Both research directions mainly utilise advanced learning-based techniques. Deep learning methods have been fostered on multiple features of pedestrians and the environment, whether derived from annotations, visual information from videos or their combinations [2]. The following two subsections are dedicated to introducing approaches that utilise DNN-based architectures for spatio-temporal analysis and feature selection/fusion.
@@ II-A Spatio-temporal Analysis
Spatio-temporal modelling can be achieved through a two-step process. Initially, visual (spatial) features per frame can be extracted using 2D convolutional neural networks (CNNs) [21] or graph convolution networks (GCNs) [16]. Subsequently, these extracted features are then fed into RNNs, such as the long short-term memory (LSTM) [15] or the gated recurrent unit (GRU) models [22, 23, 24]. For instance, in [25, 26], 2D convolutions are employed to extract visual features from image sequences, while RNNs encode the temporal relationships among these features. These sequentially encoded visual features are then inputted into a fully-connected layer to generate the ultimate intention prediction.
An alternative approach to extracting sequential visual features involves the utilisation of 3D CNNs (Conv3D) [27]. This technique directly captures spatio-temporal features by substituting the 2D kernels within the convolution and pooling layers of a 2D CNN with their 3D equivalents. For instance, in works such as [28] and [29], a framework based on a 3D CNN, specifically a 3D DenseNet, is employed to directly extract sequential visual features from sequences of pedestrian images. The ultimate prediction is then made using a fully-connected layer. Transformer architecture has also been utilised in another study [30] to tokenise the temporal input features and subsequently judge the pedestrians’ intention.
@@ II-B Feature Selection and Integration
Instead of pursuing an end-to-end approach for modelling visual features, it is possible to treat various types of information such as the pedestrian’s bounding box, body pose keypoints, vehicle movement, and the broader contextual backdrop as distinct input channels for the prediction model [12]. This necessitates the development of a fusion approach for amalgamating this diverse information.
The investigation into the types of features, such as pedestrians and environmental context, is still ongoing. For instance, pedestrian-to-vehicle distance is considered one of the most influential factors in pedestrians’ decision to cross [5]. This feature is typically estimated as a single measure between the target pedestrian and the ego vehicle [4]. Alternatively, a depth map of the scene (see Figure 2e) can be used to assess the distance from other road users and possibly reveal the underlying dynamics [31]. However, the depth map is susceptible to noise due to rough estimation, which can lead to inaccuracies in multiple pedestrian crossing scenarios [32].
Studies such as [33, 34, 35] have incorporated human poses or skeletons into pedestrian crossing prediction tasks, using pose keypoints extracted from pedestrian images to construct classifiers. This approach has shown improved prediction accuracy, but often neglects other important features or lacks attention to feature integration.
On the other hand, some studies specifically concentrated on feature integration. For instance, vision and non-vision branches fusion [36] suggest how to efficiently combine diverse data modalities at different stages of a DNN model to surge the intention prediction accuracy. Another study [10] is conducted to merge two visual and three non-visual elements of the pedestrian, scene, and subject vehicle in a multi-stream network. From a different perspective, in studies such as [11, 37], local and global contextual information has been weighted by an attention mechanism [38] and fused together to apply a prediction on Joint Attention in Autonomous Driving (JAAD) [14] and Pedestrian Intention Estimation (PIE) [15] datasets.
Figure 2: Analysing pedestrians’ features and traffic scene dynamics through: a) Pedestrian localisation, highlighted with bounding boxes, b) Gesture understanding with pose estimation, c) Object categorisation through segmentation, d) Understanding global motion patterns from optical flow, and e) Estimating distance via global depth heatmap.
@@ II-C Research Gaps
Pedestrian crossing intention highly relies on the distance of the AV to the pedestrian and the relative distance of the pedestrian to other road users which may fall into various categories of instance segmentation (e.g., cars, other pedestrians, etc.). None of the reviewed research has considered the simultaneous impact of both features on pedestrian intention.
On the other hand, to the best of our knowledge, no prior study has considered instance segmentation to smooth and normalise the distance measurement of road user instances.
In this article, we propose a new concept of Categorical Depth which integrates the classic noisy depth measurement with instance segmentation to gain more accurate depth information.
As another issue, the reviewed models, often have limited generalisability and are incapable of performing in the wild and real-world scenarios, as they have not been tested under authentic autonomous driving conditions [20]. Our study focuses on the real-world Waymo dataset which is collected from an AV’s field of view.
Lastly, there is a shortage of dedicated neural network architectures capable of effectively accommodating and extracting maximal multi-camera information from around the AV for context recognition, hence an accurate model for predicting pedestrian crossing intentions. Camera integration is proposed in this study to cope with the limited field of view.
@@ III Methodology
We propose the PIP-Net prediction model, which is based on deep neural networks for predicting pedestrian crossing intention. The model incorporates spatial-temporal features such as road users’ positioning, pose, and dynamic movements, along with a hybrid feature map that includes categorised semantic and depth information as input to the network. A multi-camera stitching and integration model is developed to facilitate panoramic viewing, enabling synchronised pedestrian ID assignment and tracking across the entire multi-view scene, thus enhancing the PIP-Net model’s understanding of spatial characteristics and contextual information.
An overview of the proposed architecture is illustrated in Figure 3. The input features are categorised into spatial kinematic data and contextual data, and they are passed to the model through distinct pipelines based on their data types. Finally, we employ recurrent and attention modules to facilitate the processing of temporal data.
Figure 3: The overview of the proposed DNN-based framework. PIP-Net receives Spatial Kinematic and Spatial Context data, and applies multi-modal feature fusion and multi-camera integration with temporal and attention-based analysis to predict pedestrians’ crossing intentions. Shift and Padding units adjust the location of the target pedestrian with respect to the corresponding camera. Aggregation module (A) fuses the cameras’ features as shown in Figure 5.
@@ III-A Spatial Kinematics
Kinematic input data includes the positioning of the pedestrian in the scene with reference to the detected pedestrian bonding box $P_{bb}$, pedestrian body pose keypoints $P_{bp}$, and the ego-vehicle speed $V_{s}$.
@@ III-B Spatial Context
Contextual input data includes pedestrian features, such as a pedestrian-bounded image (Local Content, $P_{lc}$) and corresponding motion flow analysis of the pedestrian (Local Motion, $P_{lm}$), the environment features like semantic segmentation of the scene (Semantic Context, $E_{sc}$), as well as our proposed hybrid feature map (Categorical Depth, $E_{cd}$), which refines depth information for specific pedestrians and vehicles in the scene. The mentioned features are obtained by using an ImageNet pre-trained VGG19 network as the backbone CNN, with a maximum pooling layer as suggested in [36]. Subsequently, a GRU is applied recursively to process each of these features.
@@
where $lc_{i}$ denotes the feature vector that is output by applying the CNN backbone to an RGB image. The image contains an individual pedestrian, cropped based on the bounding box location and subsequently warped to dimensions of $224\times 224\times 3$ pixels, which is the optimum spatial size in the network [30].
The resulting input feature vector is extracted as $(m,512)$ via a Conv3D layer. It is then passed through a 3D max-pooling layer (MP3D) with a kernel size of $4\times 4$ and a GRU module. This process yields a $(m,128)$ vector, where $m$ represents the observation time.
@@
where $lm_{i}$ is considered as the localised $i$-th pedestrian motion descriptor. A CNN layer is used to extract a feature vector of size $(m,256)$. This vector is then inputted into a GRU layer, resulting in an $(m,128)$ vector, suitable for concatenation with the Local Content feature vector.
@@
where $sc$ refers to the semantic segmentation of objects within the entire scene encompassing road structure and users. This feature ensures that the model considers the spatial distribution of classes for both moving and static objects within the scene. The semantic information is extracted by Slot-VPS [42] model, which is a panoptic video segmentation algorithm that not only offers semantic segments but also a unique ID for each instance of the objects in the scene. The segmented classes include 8 dynamic classes (person, rider, car, truck, bus, train, motorcycle, and bicycle) and 11 static classes (traffic light, fire hydrant, stop sign, parking meter, bench, handbag, road, sidewalk, sky, building, and vegetation).
The Categorical Depth feature $E_{cd}$ is defined as:
@@
where $cd$ represents the hybrid feature map showing the spatial distribution and distance of various instances of pedestrians and vehicles within the scene. The depth data are initially estimated using the ManyDepth model [43] and encoded in a heatmap representation, resulting in a global depth heatmap. As illustrated in Figure 2e, high-intensity spots (white and oranges) indicate proximity to the ego vehicle, while low-intensity spots (navy blue and black) represent greater distances. However, our experiments revealed that the global depth heatmap is unreliable due to inconsistencies in providing clear object boundaries. To address this, the pedestrian and vehicle instances are cropped using instance masks obtained from the Slot-VPS, as shown in Figure 4a. Subsequently, the intensity of pixels within each instance is normalised by averaging, yielding a normalised heatmap as seen in Figure 4b. This process ensures a clear and consistent depth estimation for each instance of classes, as depicted in Figure 4c.
Both inputs, $E_{sc}$ and $E_{cd}$, undergo extraction via a Conv3D layer to be assessed for the spatio-temporal analysis. The feature dimensions are gradually reduced $(512\rightarrow 256\rightarrow...\rightarrow 64)$ by repeatedly subjecting them to max-pooling layers. This process not only selects the most important information from the local neighbourhood of each pooling window but also reduces the spatial dimensions (width and height) of the feature maps and the computational complexity of the network. Then the features are organised using a flattened layer, resulting in a one-dimensional array that is suitable for concatenation and can be fed into a fully-connected layer (FC). Finally, the data is passed through a GRU module. The outputs of the three GRUs are combined and concatenated into a single output, which is then passed through an attention mechanism.
@@
Figure 4: Colour-coded visualisation of road users’ distance to the ego-vehicle after the tripe camera stitcheing. 4a is the proportion of the global depth heatmap which pedestrians and vehicles occupy; 4b is the normalised heatmap values per each instance; 4c is a visualised $E_{cd}$ feature in an RGB format in which pedestrians and vehicles integrated into the blue and red channel, respectively, and high-intensity values indicate closer users. A categorical depth map has a positive effect on removing the camera stitching effect and increasing the clarity and saliency of pedestrians and vehicles in the scene.

### v3

@@ Abstract
Accurate pedestrian intention prediction (PIP) by Autonomous Vehicles (AVs) is one of the current research challenges in this field. In this article, we introduce PIP-Net, a novel framework designed to predict pedestrian crossing intentions by AVs in real-world urban scenarios. We offer two variants of PIP-Net designed for different camera mounts and setups. Leveraging both kinematic data and spatial features from the driving scene, the proposed model employs a recurrent and temporal attention-based solution, outperforming state-of-the-art performance. To enhance the visual representation of road users and their proximity to the ego vehicle, we introduce a categorical depth feature map, combined with a local motion flow feature, providing rich insights into the scene dynamics. Additionally, we explore the impact of expanding the camera’s field of view, from one to three cameras surrounding the ego vehicle, leading to an enhancement in the model’s contextual perception. Depending on the traffic scenario and road environment, the model excels in predicting pedestrian crossing intentions up to 4 seconds in advance, which is a breakthrough in current research studies in pedestrian intention prediction. Finally, for the first time, we present the Urban-PIP dataset, a customised pedestrian intention prediction dataset, with multi-camera annotations in real-world automated driving scenarios.
@@
Autonomous vehicles, pedestrian behaviour, pedestrian crossing prediction, computer vision, deep neural networks.
@@ I Introduction
Pedestrians are the most vulnerable road users and face a high risk of fatal accidents [1]. Ensuring pedestrian safety in automated driving, particularly in mixed AV-pedestrian traffic scenarios, heavily relies on the AV’s capability in “pedestrian intention prediction (PIP)” [2]. A PIP system determines if a pedestrian is likely to cross the road shortly (within the next few seconds). This study aims to investigate the critical visual clues that pedestrians exhibit when they intend to cross the road, and then provide a model which predicts crossing behaviour, a few seconds in advance.
Anticipating pedestrian crossing behaviour is a difficult task due to various environmental factors that affect human intention [3, 4, 5]. Even in the simulated scenarios in which the majority of parameters are under control, crossing prediction is a challenging endeavour [6]. Factors like interactions with other pedestrians, traffic signs, road congestion, and vehicle speed can influence pedestrians’ tendency to cross the road in front of AVs [7].
Computer vision plays a crucial role in enabling AVs to perceive their surrounding environment by analysing the visual data captured via multiple sensors, such as cameras, LiDAR, Radar, etc. Learning-based models, in particular deep neural networks (DNNs), have shown remarkable success in various computer vision tasks, including scene understanding, semantic segmentation [8], road users classification, localisation [9], and motion prediction [10]. Figure 1 illustrates some of the perceivable factors such as depth, pedestrian pose, and surrounding objects, that an autonomous vehicle should consider to interpret the scene and estimate the pedestrians’ intention. DNNs are particularly effective at learning complex patterns and features from visual data, making them a natural fit for tasks that involve analysing images or videos to comprehend pedestrian behaviour [11, 12, 13]. They also offer significant capabilities in multi-modal integration by providing a neural-based mechanism to process and fuse all the perceived information from diverse sensors. This integration may enhance the overall understanding of the environment and help to make more accurate and safer decisions [14].
@@
Several datasets, such as JAAD [15], PIE [16], and STIP [17], use onboard camera recordings and their data are publicly released for the study of pedestrians’ behaviour before and during road crossing. However, most of the current research works suffer from supplying a multi-camera setup to leverage the benefits of sensor fusion and multi-modal perception. In addition to the above-mentioned datasets, some baseline approaches [18] have also been established for analysing the visual cues and signals that pedestrians emit through their body language and positioning. The approaches highlight the benefits of combining these features with contextual information. Contextual information may include factors such as the road’s location, the time of day, weather conditions, the presence of traffic signals or crosswalks, the type of road (urban, suburban, rural), and the position and behaviour of other vehicles near the scene [19]. To the best of our knowledge, no extensive research has been conducted to understand and interpret such contextual details and their effects on pedestrian’s decision-making.
In this study, we propose a customised DNN-based framework, called “PIP-Net” that takes various features of pedestrians, the environment, and the ego-vehicle state into account, to learn the context of a crossing scenario and consequently predict the crossing intention of pedestrians in real-world AV urban driving scenarios. The main contributions of this research are highlighted as follows:
@@
  • Finally, we evaluate the effectiveness of the proposed model on the widely utilised PIE dataset and the introduced Urban-PIP dataset, outperforming the state-of-the-art (SOTA) for crossing prediction.
@@ II Related Works
Recently, pedestrian crossing intention prediction research has surged and gained significant attention within the autonomous driving research community [11, 12, 13]. Most current methods mainly address the problem by taking two aspects into account: Discovering influential factors and features for interpreting road users’ interactions [6, 7, 4], and designing the analytic model to predict the pedestrians’ crossing intention [19, 15, 16, 17, 18]. Both research directions mainly utilise advanced learning-based techniques. Deep learning methods have been fostered on multiple features of pedestrians and the environment, whether derived from annotations, visual information from videos or their combinations [2, 21]. The following two subsections are dedicated to introducing approaches that utilise DNN-based architectures for spatio-temporal analysis and feature selection/fusion.
@@ II-A Spatio-temporal Analysis
Spatio-temporal modelling can be achieved through a two-step process. Initially, visual (spatial) features per frame can be extracted using a visual encoder such as 2D convolutional neural networks (CNNs) [22], vision transformers (ViTs) [23], or graph convolution networks (GCNs) [17]. Subsequently, these extracted features are then fed into a temporal encoder like recurrent neural networks (RNNs) [24], using long short-term memory (LSTM) [25, 26, 27] or the gated recurrent unit (GRU) modules [28, 29, 30]. For instance, in [26, 31, 24], 2D convolutions are employed to extract visual features from image sequences, while RNNs encode the temporal relationships among these features. These sequentially encoded visual features are then inputted into a fully-connected layer to generate the ultimate crossing prediction.
An alternative approach is extracting both spatial and temporal features involves the utilisation of 3D CNNs (Conv3D) [32] or Transformer architecture [33]. 3D CNNs can directly capture spatio-temporal features by substituting the 2D kernels within the convolution and pooling layers of a 2D CNN with their 3D equivalents. For instance, in works such as [34, 35, 36, 37], a framework based on a 3D CNN is employed to directly extract sequential visual features from sequences of pedestrian images. The ultimate prediction is then made using a fully-connected layer. Transformer architecture uses self-attention mechanisms [38] to capture long-range relationships, both within a single frame (spatial) and between different frames (temporal). This helps in thoroughly analysing pedestrian dynamics [39, 40, 41, 42]. Furthermore, hybrid models combining CNNs, RNNs, and attention mechanisms have been explored [43, 44, 45] to leverage the strengths of these approaches.
@@ II-B Feature Selection and Integration
Reliance on a single pedestrian feature for crossing prediction, such as using only the pedestrian’s pose kinematics [46] or bounding box [41], has resulted in context-free crossing predictions that neglect other scene-specific modalities and miss traffic and situational awareness. Instead, it is possible to treat various types of information, such as the pedestrian’s image, body pose keypoints, bounding box, vehicle dynamics, and the broader contextual backdrop as distinct input channels for the prediction model [13].
Studies such as [47, 48, 49] have incorporated human poses or skeletons into pedestrian crossing prediction tasks alongside various features such as pedestrian’s local image, bounding box, and vehicle speed, to construct the intention classifier. This approach has shown improved prediction accuracy but often neglects other important features or lacks proper attention to feature integration.
The investigation into the types of features, such as pedestrians and environmental context, is still ongoing. For instance, pedestrian-to-vehicle distance is considered one of the most influential factors in pedestrians’ decisions to cross [6]. This feature is typically estimated as a single measure between the target pedestrian and the ego vehicle [4, 50], while relative velocity and closing speed, which refers to how quickly the gap between the ego vehicle and the pedestrian is decreasing, are also investigated [51]. Alternatively, a depth map of the scene (see Figure 2e) can be used to assess the distance from other road users and possibly reveal the underlying dynamics [52]. However, the depth map is susceptible to noise due to rough estimation, which can lead to inaccuracies in scenarios involving multiple pedestrians crossing [53].
On the other hand, some studies specifically concentrated on feature integration. For instance, vision and non-vision branches fusion [43] suggest how to efficiently combine diverse data modalities at different stages of a DNN model to surge the intention prediction accuracy. Another study [11] is conducted to merge two visual and three non-visual elements of the pedestrian, scene, and subject vehicle in a multi-stream network. From a different perspective, in studies such as [12, 45], local and global contextual information has been weighted by an attention mechanism and fused together to apply a prediction on Joint Attention in Autonomous Driving (JAAD) [15] and Pedestrian Intention Estimation (PIE) [16] datasets.
@@ II-C Research Gaps
Despite advancements in pedestrian crossing intention prediction, several key gaps persist, as outlined in recent surveys [2, 21, 5]. In this study, we focus on addressing four research gaps:
1) Pedestrian crossing intention highly relies on the distance of the AV to the pedestrian and the relative distance of the pedestrian to other road users, which may fall into various categories of instance segmentation (e.g., cars, other pedestrians, etc.). None of the reviewed research has considered the simultaneous impact of both features on pedestrian intention.
2) To the best of our knowledge, no prior study has considered instance segmentation to smooth and normalise the distance measurement of road user instances. In this article, we propose a new concept of Categorical Depth which integrates the classic noisy depth measurement with instance segmentation to gain more accurate depth information and provide context-aware spatial data.
3) The reviewed models, often have limited generalisability and are incapable of performing in the wild and real-world automated driving scenarios, as they have normally been tested with a human-driven vehicle [54]. Our study focuses on the real-world Waymo dataset, which is collected from an AV’s field of view. Our added pedestrian intention annotations to this dataset help address the issue of inadequate annotated ground-truth data availability in this field of research.
4) There is a shortage of dedicated neural network architectures capable of effectively accommodating and extracting maximal multi-camera information from around the AV for context recognition, hence an accurate model for predicting pedestrian crossing intentions. Camera integration is proposed in this study to cope with the limited field of view.
Figure 2: Analysing pedestrians’ features and traffic scene dynamics through: a) Pedestrian localisation, highlighted with bounding boxes, b) Gesture understanding with pose estimation, c) Object categorisation through segmentation, d) Global motion patterns using optical flow, and e) Estimating distance via a global depth heatmap.
@@ III Methodology
We propose the PIP-Net model, which is based on deep neural networks for predicting pedestrian crossing intention. The model incorporates spatial-temporal features such as road users’ positioning, pose, and dynamic movements, along with a hybrid feature map, the categorised semantic and depth information as input to the network. A multi-camera stitching and integration model is developed to facilitate panoramic viewing, enabling a synchronised pedestrian ID assignment and tracking across the entire multi-view scene, thus enhancing the PIP-Net model’s understanding of spatial characteristics and contextual information.
An overview of the proposed architecture is illustrated in Figure 3. The input features are categorised into spatial kinematic data and contextual data, and they are passed to the model through distinct pipelines based on their data type. Finally, the recurrent module and attention module are utilised to improve temporal data processing.
Figure 3: The overview of the proposed DNN-based framework. PIP-Net receives Spatial Kinematic and Spatial Context data, and applies multi-modal feature fusion and multi-camera integration with temporal and attention-based analysis to predict the pedestrians’ crossing intentions. Shift and Padding units adjust the location of the target pedestrian with respect to the corresponding camera. Aggregation module (A) fuses the cameras’ features as shown in Figure 5.
@@ III-A Spatial Kinematics
Kinematic input data includes the positioning of the pedestrian in the scene with reference to the detected pedestrian bounding box $P_{bb}$, pedestrian body pose keypoints $P_{bp}$, and the ego-vehicle speed $V_{s}$.
@@ III-B Spatial Context
Contextual input data includes pedestrian features, such as a pedestrian-bounded image (Local Content, $P_{lc}$) and corresponding motion flow analysis of the pedestrian (Local Motion, $P_{lm}$), the environment features like semantic segmentation of the scene (Semantic Context, $E_{sc}$), as well as our proposed hybrid feature map (Categorical Depth, $E_{cd}$), which refines depth information for specific pedestrians and vehicles in the scene.
@@
where $lc_{i}$ denotes the feature vector that is output by applying the CNN backbone to an RGB image. An ImageNet pre-trained VGG19 network is used as the CNN backbone, with a maximum pooling layer as suggested in [43]. Subsequently, a GRU is applied recursively to process the temporal dimesion of the feature. The RGB image contains an individual pedestrian, cropped based on the bounding box location and subsequently warped to dimensions of $224\times 224\times 3$ pixels, which is the optimum spatial size in the network [23].
@@
where $lm_{i}$ is considered as the localised $i$-th pedestrian motion descriptor. A Conv3D layer is used to extract a feature vector of size $(m,512)$, where $m$ represents the observation sequence length. The fueature vector is then passed through a 3D max-pooling layer (MP3D) with a kernel size of $4\times 4$ and a GRU module. This process yields a $(m,128)$ vector, this vector is then inputted into a GRU layer, resulting in an $(m,128)$ vector, suitable for concatenation with the Local Content feature vector.
@@
where $sc$ refers to the semantic segmentation of objects within the entire scene encompassing road structure and users. This feature ensures that the model considers the spatial distribution of classes for both moving and static objects within the scene. The semantic information is extracted by Slot-VPS [58] model, which is a panoptic video segmentation algorithm that not only offers semantic segments but also a unique ID for each instance of the objects in the scene. The segmented classes include 8 dynamic classes (person, rider, car, truck, bus, train, motorcycle, and bicycle) and 11 static classes (traffic light, fire hydrant, stop sign, parking meter, bench, handbag, road, sidewalk, sky, building, and vegetation). The instance segments corresponding to pedestrians and vehicles are integrated into the Categorical Depth feature, denoted as $E_{cd}$, using the following formulation:
@@
where $cd$ represents the hybrid feature map containing depth information for pedestrians and vehicles within the scene. The depth data are initially estimated using the ManyDepth model [59] and encoded in a heatmap representation, resulting in a global depth heatmap. As illustrated in Figure 2e, high-intensity spots (white and oranges) indicate proximity to the ego vehicle, while low-intensity spots (navy blue and black) represent greater distances. However, our experiments revealed that the global depth heatmap is unreliable due to inconsistencies in providing clear object boundaries. To address this, the pedestrian and vehicle instances are cropped using instance masks obtained from the Slot-VPS, as shown in Figure 4a. Subsequently, the intensity of pixels within each instance is normalised by averaging, yielding a normalised heatmap as seen in Figure 4b. This process ensures a clear and consistent depth estimation for each class, as depicted in Figure 4c.
@@
Figure 4: Colour-coded visualisation of road users’ distance to the ego-vehicle after the tripe camera stitching. 4a is the proportion of the global depth heatmap which pedestrians and vehicles occupy; 4b is the normalised heatmap values per each instance; 4c is a visualised $E_{cd}$ feature in an RGB format in which pedestrians and vehicles integrated into the blue and red channel, respectively, and high-intensity values indicate closer users. A categorical depth map has a positive effect on removing the camera stitching effect and increasing the clarity and saliency of pedestrians and vehicles in the scene.
Both inputs, $E_{sc}$ and $E_{cd}$, undergo extraction via three Conv3D layers to be assessed for the spatio-temporal analysis. The feature dimensions are gradually reduced $(512\rightarrow 256\rightarrow 128\rightarrow 64)$ by repeatedly subjecting them to max-pooling layers. This process not only selects the most important information from the local neighbourhood of each pooling window but also reduces the spatial dimensions (width and height) of the feature maps and the computational complexity of the network. Then the features are organised using a flattened layer, resulting in a one-dimensional array that is suitable for concatenation and can be fed into a fully-connected layer (FC). Finally, the data is passed through a GRU module. The outputs of the three GRUs are combined and concatenated into a single output, which is then passed through an attention mechanism.